Skip to Search

 
#

Support Services Meeting Notes
July 14, 2008

Present: Bill Covey, Will Chaney, Lori Driscoll, Erich Kesse, Rob Roberts
Facilities
  1. Staff chairs for Systems - ten are needed. Folding chairs will work. They are currently using extras from Room 1 A.
  2. West furniture rearrangement - finished two floors. Domes installed on the round tables. The table anchors are here; we are painting them.
  3. Office moves underway for Directors: Sam Huang and Alicia Antone.
  4. MSL - not moving yet. We did one office move this morning.
  5. Office and counter moves in Reference - soon. Cabinet has been ordered.
  6. August 9-12 - escalator cleaning scheduled. West will be closed for two of those days.
  7. Door construction to start at MSL today.
  8. Antique display case to be moved by Turner. Should it go to Special Collections or 3rd floor West? Rob to verify with Shelley.
Access Support
  1. Lori, Bill Merrie Davidson, Alicia Antone, Nancy Poehlmann and Jay Smith met with the Institute of Black Culture. They are interested in running a Circulation system for the check-out of their books. We reviewed the needs for security, patron policy, etc. Their inventory is 1000 items or less. We discussed alternative systems. Lori will help them set up policies and procedures. Bill spoke with Winston about a template for access. Perhaps a simple application to track the status on the Web. They may go that route. They really don't want to be burdened with a whole Circ system.
Systems
  1. We have one chair in the conference room that needs to be repaired.
  2. The high-end scanners bought for the Library are going out of warranty. The company does not want to support the repairs. They suggested we buy new scanners. Bill would like Will to send him the contact information and details for the company.
  3. UFirt - required a patch on DNS server. Ours are patch, Bill notes the CNS DNS server needed to patch at the same time.
  4. Disk corruption on the new SAN file server. We failed over to the backup server.
DLC
  1. Erich met with John Ingram, Carl VanNess, Flo Turcotte, and John Nemmers about the Jeremie Collection. John Nemmers decided to back off and talk with Sam Huang for outside support.
  2. The JPEG2K server is stuffed to the gills. 1.4 TB is 97% full. Erich passed out a spreadsheet with numbers showing the projected rate of increase. Discussions of the specs. Discussions leads into talk of what DLC servers are backed up. Bill would like this information. Erich wants to look into buying more storage space with grants. There is money in the Bryant project to buy more space for the JPEG2K server.
Directors
  1. The Directors had a meeting on various aspects of HR. Discussed training institutes. There will be very limited slots (6 or so) for development.
  2. LibQual - Climate Qual we will pass on.
  3. Discussion on the status of searches.
  4. Various budget discussions. No word on IT equipment budget.
  5. Reorg - no blog commentary.
  6. AD tech services position - resumes to be looked at on 7/14.

 

University of Florida Home Page